OpenStudy (anonymous):

a bus covers a distance of 60 km in 1 1/4 hr.how far would it travel in 1 1/2 hr

OpenStudy (anonymous):

@zepdrix

OpenStudy (jdoe0001):

\(\large \begin{array}{ccllll} km&hr \\\hline\\ 60&1\frac{1}{4}\\ x&1\frac{1}{2} \end{array}\implies \cfrac{60}{x}=\cfrac{1\frac{1}{4}}{1\frac{1}{2}}\)

OpenStudy (anonymous):

@Kainui

OpenStudy (anonymous):

a: 70 b. 50 c. 72 d. 80 didn't getting anyone by your method ? @jdoe0001

OpenStudy (jdoe0001):

hmm

OpenStudy (jdoe0001):

what did you get ?

OpenStudy (anonymous):

120

OpenStudy (jdoe0001):

one sec

OpenStudy (jdoe0001):

\(\bf \cfrac{60}{x}=\cfrac{1\frac{1}{4}}{1\frac{1}{2}}\implies \cfrac{60}{x}=\cfrac{\frac{5}{4}}{\frac{3}{{\color{blue}{ 2}}}}\implies \cfrac{60}{x}=\cfrac{5}{4}\times \cfrac{{\color{blue}{ 2}}}{3} \\ \quad \\ \implies \cfrac{60}{x}=\cfrac{5\times \cancel{ 2 }}{\cancel{ 4 }\times 3}\implies \cfrac{60}{x}=\cfrac{5}{6}\)

OpenStudy (jdoe0001):

what would that give you?

OpenStudy (anonymous):

72 and that's the answer . :) Thank you so much ^_^. actually I'm not a maths student , to pass a test I'm practicing general maths So I'm newbie :P

OpenStudy (jdoe0001):

yw